OpenStudy (anonymous):

fast food restaurant sells 425 large orders of fries per week at price $2.75 each. A market survey indicates that for each $0.10 decrease in price, sales will increase by 20 orders of fries. What is the demand function. I tried doing it, but the way I was told to do it didn't give me the right answer

OpenStudy (anonymous):

You're told information to find slope; our demand function is written \(p=q\) and we're told that for \(\Delta p=-0.1\) we have \(\Delta q=20\) so our slope is:$$\frac{\Delta p}{\Delta q}=-0.005$$

OpenStudy (anonymous):

We're also told a point on your curve, namely \(p=2.75,q=245\). Presuming a linear demand function, we have \(p=a-bq=a-0.005q\). Plug in and find \(a\).

OpenStudy (anonymous):

can I show you the way I did it

OpenStudy (anonymous):

let p= price of fries let n=# of 0.10 decrease let x=# of fries sold p=2.75-0.10n x=425+20n then I solve for n in the x function and sub it into the p function

OpenStudy (anonymous):

but i don't get the books answer